### v2

@@ 4 Conditional OT-Flow (COT-Flow)
In this section, we extend the OT-regularized continuous normalizing flows in [40] to the conditional generative modeling problem introduced in Section 2.
@@
Following the general approach of continuous normalizing flows [10, 22], we express the transport map that pushes forward the reference distribution $\rho_{{\bf z}_{x}}({\bf z}_{x})$ to the target $\pi({\bf x}|{\bf y})$ via the flow map of an ODE. That is, we define $g({\bf z}_{x};{\bf y})={\bf u}(1)$ as the terminal state of the ODE
@@
Here, the terminal and running costs in the objective functional are the $L_{2}$ anti-derivatives of $J_{\rm NLL}$ and $P_{\rm DOT}$, respectively, and the continuity equation is used to represent the density evolution under the velocity $v$. To be precise, (15) can be seen as a discretization of (18) in the Lagrangian coordinates defined by the reversed ODE (26). Note that both formulations differ in the way the transport costs are measured: (16) computes the cost of pushing the conditional distribution to the Gaussian, while (18) penalizes the costs of pushing the Gaussian to the conditional. While these two terms do not generally agree, they coincide for the $L_{2}$ optimal transport map; see [19, Corollary 2.5.13] for a proof for unconditional optimal transport. For more insights into MFGs and their relation to optimal transport and generative modeling, we refer to our prior work [50] and the more recent work [62, Sec 3.4]. The fact that the solutions of the microscopic version in (15) and the macroscopic version in (18) agree is remarkable and was first shown in the seminal work [31].
@@
(26) $\begin{split}\min_{\theta}&\quad J_{\rm NLL}[g_{\theta}]+\alpha_{1}P_{\rm DOT}[\nabla_{\bf p}\Phi_{\theta}]+\alpha_{2}P_{\rm HJB}[\Phi_{\theta}]\\ \text{ where }&\quad g_{\theta}^{-1}({\bf x};{\bf y})={\bf p}(0)\;\text{and}\;\frac{d}{dt}{\bf p}(t)=-\frac{1}{\alpha_{1}}\nabla_{\bf p}\Phi_{\theta}(t,{\bf p}(t);{\bf y}),\;t\in(0,1],\;{\bf p}(1)={\bf x}.\end{split}$
@@
After training, we draw i.i.d. samples from the approximated conditional distribution for a given ${\bf y}$ by sampling ${\bf z}_{x}\sim\rho_{{\bf z}_{x}}({\bf z}_{x})$ and solving the ODE (13). Since we train the inverse generator using a fixed number of integration steps, $n_{t}$, it is interesting to investigate how close our solution is to the continuous problem. One indicator is to compare the samples obtained with different number of integration steps during sampling. Another indicator is to compute the variance of $\nabla\Phi$ along the trajectories.
@@
As described in Section 5, we randomly sample the width, $w$, of the neural network, the number of time integration steps during training, the penalty parameters $\alpha_{1},\alpha_{2}$, and the hyperparameters of the Adam optimizer (batch size and initial learning rate) from the values listed in Table 2.
@@ 5 Implementation and Experimental Setup
This section describes our implementations, experimental setups and provides guidance for applying our techniques to new problems.
@@
The scripts for implementing our neural network approaches and running our numerical experiments are written in Python using PyTorch. For datasets that are not publicly available, we provide the binary files we use in our experiments and the Python scripts for generating the data. We have published the code and data along with detailed instructions on how to reproduce the results in our main repository https://github.com/EmoryMLIP/PCP-Map.git. Since the COT-Flow approach is a generalization of a previous approach, we have created a fork for this paper https://github.com/EmoryMLIP/COT-Flow.git.
@@
Finding hyperparameters, such as network architectures and optimization parameters, is crucial in neural network approaches. In our experience, the choice of hyperparameters is often approach and problem-dependent. Establishing rigorous mathematical principles for choosing the parameters in these models (as is now well known for regularizing convex optimization problems based on results from high-dimensional statistics [39]) is an important area of future work and is beyond the scope of this paper. Nevertheless, we present an objective and robust way, relying only on random grid search over a pre-determined search space Table 2, of identifying an effective combination of hyperparameters for our approaches. To be more specific, we first perform a pilot run for each experiment using randomly selected hyperparameters combinations from a subset of Table 2. The number of hyperparameters considered in the pilot run for each experiment will be presented in the next section. Then, after identifying the best combination in terms of validation loss, we perform full training with that combination.
Table 2: Hyperparameter search space for PCP-Map and COT-Flow. Here $m$ denotes the size of the context feature or observation ${\bf y}$
@@
[…]
@@ 6 Numerical Experiments
We test the accuracy, robustness, efficiency, and scalability of our approaches from Sections 3 and 4 using three problem settings that lead to different challenges and benchmark methods for comparison. All numerical expriments are conducted on one Quadro RTX 8000 GPU with 48 GB of RAM. In Section 6.1, we compare our proposed approaches to the Adaptive Transport Maps (ATM) approach developed in [3] on estimating the joint and conditional distributions of six UCI tabular datasets [32]. In Section 6.2, we compare our approaches to a provably convergent approximate Bayesian computation (ABC) approach on accuracy and computational cost using the stochastic Lotka–Volterra model, which yields an intractable likelihood. Using this dataset, we also compare PCP-Map’s and COT-Flow’s sampling efficiency for different settings. In Section 6.3, we demonstrate the scalability of our approaches to higher-dimensional problems by comparing them to the flow-based neural posterior estimation (NPE) approach on an inference problem involving the 1D shallow water equations. To demonstrate the improvements of PCP-Map over the amortized CP-Flow approach in the repository associated with  [24], we compare computational cost in Section 6.4.
@@ 6.1 UCI Tabular Datasets
(30) $J_{{\rm NLL}}[T]={\rm\mathbb{E}}_{\pi({\bf x},{\bf y})}\left[{\frac{1}{2}}\left\|T^{-1}({\bf x},{\bf y})\right\|^{2}-\log\det\nabla T^{-1}({\bf x},{\bf y})\right].$
@@
(31) $J_{{\rm NLL},{\bf y}}[\mathsf{h}]=\mathbb{E}_{{\bf y}\sim\pi({\bf y})}\left[\left\|\mathsf{h}^{-1}({\bf y})\right\|^{2}-\log\det\nabla_{\bf y}\mathsf{h}^{-1}({\bf y})\right].$
@@
for $k=0,\ldots,K-1$. Here $\sigma^{(s)}$ is the softplus activation function. Since the FICNN map is not part of our contribution, we use the input-augmented ICNN implementation used in [24]. For the second map component $g^{-1}$, we use the gradient of a PICNN as described in Section 3. For COT-Flow, we construct two distinct neural network parameterized potentials $\Phi_{x}$ and $\Phi_{y}$. Here, $\Phi_{y}$ only takes the conditioning variable ${\bf y}$ as inputs and can be constructed exactly like in [40]. $\Phi_{x}$ acts on all features and is the same potential described in Section 4. We learn the inverse conditional generator $g^{-1}_{\theta}$ to perform conditional density estimation. This is equivalent to solely learning the weights $\theta$ of the PICNN $\tilde{G}_{\theta}$ for PCP-Map. For COT-Flow, we only construct and learn the potential $\Phi_{x}$.
Table 3: Hyperparameter sample space for the UCI tabular datasets experiment.
@@
The hyperparameter sample space we use for this experiment is presented in Table 3. We select smaller batch sizes and large learning rates as this leads to fast convergence on these relatively simple problems. For each dataset, we select the ten best hyperparameter combinations based on a pilot run for full training. For PCP-Map’s pilot run, we performed 15 epochs for all three conditional datasets, three epochs for the Parkinson’s and the White Wine datasets, and four epochs for Red Wine using 100 randomly sampled combinations. For COT-Flow, we limit the pilot runs to only 50 sampled combinations due to a narrower sample space on model architecture. To assess the robustness of our approaches, we performed five full training runs with random initializations of network weights for each of the ten hyperparameter combinations for each dataset.
In Table 4, we report the best, median, and worst mean negative log-likelihood on the test data across the six datasets for the two proposed approaches and the best results for ATM. The table demonstrates that the best models outperform ATM for all datasets and that the median performance is typically superior. Overall, COT-Flow slightly outperforms PCP-Map in terms of loss values for the best models. The improvements are more pronounced for the conditional sampling tasks, where even the worst hyperparameters from PCP-Map and COT-Flow improve over ATM with a substantial margin.
Table 4: Mean negative log-likelihood comparisons between ATM, PCP-Map, and COT-Flow on test data. For our approaches, we report the best, median, and worst results over different hyperparameter combinations and five training runs. Lower is better, and we highlight the best results in bold.
@@
6.2 Stochastic Lotka-Volterra
We compare our approaches to an ABC approach based on Sequential Monte Carlo (SMC) for likelihood-free Bayesian inference using the stochastic Lotka-Volterra (LV) model [60]. The LV model is a stochastic process whose dynamics describe the evolution of the populations $S(t)=(S_{1}(t),S_{2}(t))$ of two interacting species, e.g., predators and prey. These populations start from a fixed initial condition $S(0)=(50,100)$. The parameter ${\bf x}\in\mathbb{R}^{4}$ determines the rate of change of the populations over time, and the observation ${\bf y}\in\mathbb{R}^{9}$ contains summary statistics of the time series generated by the model. This results in an observation vector with nine entries: the mean, the log-variance, the auto-correlation with lags one and two, and the cross-correlation coefficient. The procedure for sampling a trajectory of the species populations is known as Gillespie’s algorithm.
@@ 6.2 Stochastic Lotka-Volterra
Figure 2: Posterior samples in log scale and MAP point quality comparisons between proposed approaches and ABC with ${\bf x}^{*}=(0.01,0.5,1,0.01)^{\top}$. Left: posterior samples generated by proposed approaches trained on 50k samples. Middle: posterior samples generated by proposed approaches trained on 500k samples. The red dots and bars correspond to ${\bf x}^{*}$, and the black crosses and bars correspond to the MAP point. Right: posterior samples from ABC.
@@
We evaluate our approaches for maximum-a-posteriori (MAP) estimation and posterior sampling. We consider a true parameter ${\bf x}^{*}=(0.01,0.5,1,0.01)^{\top}$, which was chosen to give rise to oscillatory behavior in the population time series. Given one observation ${\bf y}^{*}\sim\pi({\bf y}|{\bf x}^{*})$, we first identify the MAP point by maximizing the estimated log-likelihoods provided by our approaches. Then, we generate 2000 samples from the approximate posterior $\pi({\bf x}|{\bf y}^{*})$ using our approaches. Fig. 2 presents one and two-dimensional marginal histograms and scatter plots of the MAP point and samples, compared against 2000 samples generated by the SMC-ABC algorithm from [5]. Our approaches yield samples tightly concentrated around the MAP points that are close to the true parameter ${\bf x}^{*}$.
To provide more evidence that our learning approaches indeed solve the amortized problem, Fig. 3 shows the MAP points and approximate posterior samples generated from a new random observation ${\bf y}^{*}$ corresponding to the true parameter ${\bf x}^{*}=(0.02,0.02,0.02,0.02)^{\top}$. We observe similar concentrations of the MAP point and posterior samples around the true parameter and similar correlations learned by the generative model and ABC, for example, between the third and other parameters.
@@
Figure 3: Posterior samples log scale and MAP point quality comparisons between proposed approaches and ABC with ${\bf x}^{*}=(0.02,0.02,0.02,0.02)^{\top}$. Left: posterior samples generated by proposed approaches trained on 50k samples. Middle: posterior samples generated by proposed approaches trained on 500k samples. The red dots and bars correspond to ${\bf x}^{*}$, and the black crosses and bars correspond to the MAP point. Right: posterior samples from ABC.
To further validate that we approximate the posterior distribution as well, Fig. 4 compares the population time series generated from approximate posterior samples given by our approaches and SMC-ABC. This corresponds to comparing the posterior predictive distributions for the states $S(t)$ given different posterior approximations for the parameters. While the simulations have inherent stochasticity, we observe that the generated posterior parameter samples from all approaches recover the expected oscillatory time series simulated from the true parameter ${\bf x}^{\ast}$, especially at earlier times.
Figure 4: Posterior predictives quality comparisons between proposed approaches and ABC with ${\bf x}^{*}=(0.01,0.5,1,0.01)^{\top}$. The solid lines in each plot represent a simulated time series using ${\bf x}^{*}$. Dotted lines represent time series simulated using ten randomly selected posterior samples from the 2000. Left: posterior predictives from proposed approaches trained on 50k samples. Middle: posterior predictives from proposed approaches trained on 500k samples. Right: posterior predictives from ABC.
In the experiments above, we employed $n_{t}=32$ to generate posterior samples during testing for the COT-Flow. However, one can also decrease $n_{t}$ after training to generate samples faster without sacrificing much accuracy, as shown in Fig. 5. Thereby, one can achieve faster sampling speed using COT-Flow than using PCP-Map as demonstrated in Table 5. To establish such a comparison, we increase the $l$-BFGS tolerance when sampling using PCP-Map, which produces a similar effect on sampling accuracy than decreasing the number of time steps for COT-Flow.
Figure 5: Relative normed errors between COT-Flow posterior samples generated with smaller $n_{t}$ values and samples generated with $n_{t}=32$. Here ${\bf x}^{*}_{1}=(0.01,0.5,1,0.01)^{\top}$ and ${\bf x}^{*}_{2}=(0.02,0.02,0.02,0.02)^{\top}$.
Table 5: Sampling efficiency comparisons between PCP-Map and COT-Flow in terms of GPU time in seconds (s). We report the mean and standard deviation over five runs, respectively.
@@
The shallow water equations model wave propagation through shallow basins described by the depth profiles parameter, ${\bf x}\in\mathbb{R}^{100}$, which is discretized at 100 equidistant points in space. After solving the equations over a time grid with 100 cells, the resulting wave amplitudes form the 10k-dimensional raw observations. As in [48], we perform a 2D Fourier transform on the raw observations and concatenate the real and imaginary parts since the waves are periodic. We define this simulation-then-Fourier-transform process as our forward model and denote it as $\Psi({\bf x})$. Additive Gaussian noise is then introduced to the outputs of $\Psi$, which gives us the observations ${\bf y}=\Psi({\bf x})+0.25{\boldsymbol{\epsilon}}$, where ${\bf y},{\boldsymbol{\epsilon}}\in\mathbb{R}^{200\times 100}$ and ${\boldsymbol{\epsilon}}_{i,j}\sim\mathcal{N}(0,1)$. We aim to use the proposed approaches to learn the posterior $\pi({\bf x}|{\bf y})$.
@@
Table 6: Hyperparameter sample space for the 1D shallow water equations experiment.
@@
$[\log(\alpha_{1}),\log(\alpha_{2})]$ [$\mathcal{U}(10^{2},10^{5})$, $\mathcal{U}(10^{2},10^{5})$]

### v3

@@ 4 Conditional OT-Flow (COT-Flow)
In this section, we approximate the COT map $g$ with a continuous normalizing flow (CNF)  [10, 22] whose time-dependent velocity field is the partial gradient of a scalar-valued neural network. Similarly to the previous section, we learn the parameters of the network through maximum likelihood training but add an OT regularization term to enforce uniqueness and favor simple generators with straight trajectories. To this end, we extend the OT-regularized CNF in [41] to the conditional setting.
@@
Following the general approach of CNFs, we represent a $g$ that pushes forward the reference distribution $\rho_{{\bf z}_{x}}({\bf z}_{x})$ to the target distribution $\pi({\bf x}|{\bf y})$ via the flow map of an ODE [38]. That is, we define $g({\bf z}_{x};{\bf y})={\bf u}(1)$ as the terminal state of the ODE
@@
Here, the terminal and running costs in the objective functional are the $L_{2}$ anti-derivatives of $J_{\rm NLL}$ and $P_{\rm DOT}$, respectively, and the continuity equation is used to represent the density evolution under the velocity $v$. To be precise, (15) can be seen as a discretization of (18) in the Lagrangian coordinates defined by the reversed ODE (26). Note that both formulations differ in the way transport costs are measured: (16) computes the cost of pushing the conditional distribution to the Gaussian, while (18) penalizes the costs of pushing the Gaussian to the conditional. While these two terms do not generally agree, they coincide for the $L_{2}$ optimal transport map; see [19, Corollary 2.5.13] for a proof for unconditional optimal transport. For more insights into MFGs and their relation to optimal transport and generative modeling, we refer to our previous work [52] and the more recent work [65, Sec 3.4]. The fact that the solutions of the microscopic version in (15) and the macroscopic version in (18) agree is remarkable and was first shown in the seminal work [33].
@@
(26) $\begin{split}\min_{\theta}&\quad J_{\rm NLL}[g_{\theta}^{-1}]+\alpha_{1}P_{\rm DOT}[\nabla_{\bf p}\Phi_{\theta}]+\alpha_{2}P_{\rm HJB}[\Phi_{\theta}]\\ \text{ where }&\quad g_{\theta}^{-1}({\bf x};{\bf y})={\bf p}(0)\;\text{and}\;\frac{d}{dt}{\bf p}(t)=-\frac{1}{\alpha_{1}}\nabla_{\bf p}\Phi_{\theta}(t,{\bf p}(t);{\bf y}),\;t\in(0,1],\;{\bf p}(1)={\bf x}.\end{split}$
@@
After training, we draw i.i.d. samples from the approximate conditional distribution for a given ${\bf y}$ by sampling ${\bf z}_{x}\sim\rho_{{\bf z}_{x}}({\bf z}_{x})$ and solving the ODE (13). Since we train the inverse generator using a fixed number of integration steps, $n_{t}$, it is interesting to investigate how close our solution is to the continuous problem. One indicator is to compare the samples obtained with different number of integration steps during sampling. Another indicator is the variance of $\nabla\Phi$ along the trajectories.
@@
As described in Section 5, we select the width $w$ of the neural network, the number of time integration steps during training, the penalty parameters $\alpha_{1},\alpha_{2}$, and the hyperparameters of the Adam optimizer (batch size and initial learning rate) from the values listed in Table 2. Following the original implementation of OT-Flow [41], we fix the network depth $K$ to 2. For more complex problems, we recommend exploring wider, deeper networks or more advanced architectures.
@@ 5 Implementation and Experimental Setup
This section describes our implementations, experimental setups, and provides guidance for applying our techniques to new problems.
@@
The scripts for implementing our neural network approaches and running our numerical experiments are written in Python using PyTorch. For datasets that are not publicly available, we provide the binary files we use in our experiments and the Python scripts for generating the data. We have published the code and data along with detailed instructions on how to reproduce the results in our main repository, https://github.com/EmoryMLIP/PCP-Map.git. Since COT-Flow is a generalization of a previous approach, we have created a fork for this paper at https://github.com/EmoryMLIP/COT-Flow.git.
@@
Setting the values of key hyperparameters, describing both the network architecture and the optimization (training) algorithm, is crucial for achieving good performance from neural network approaches. The best values of such hyperparameters are typically both algorithm- and problem-specific. Ideally, one would like to have rigorous mathematical principles for choosing hyperparameters (e.g., akin to well-established methods for regularizing convex optimization problems in high-dimensional statistics [40]), but this remains a largely open problem, outside the present scope. Usual practice is instead to tune hyperparameters via some black-box optimization method, such as Bayesian optimization [53].
Rather than taking such an approach here, we use a simpler optimization approach that also better demonstrates the robustness of our algorithms to sub-optimal hyperparameter values. We begin with the tensor product of candidate hyperparameter values given in Table 2 (with problem-specific modifications reported in Table 3 and Table 6). From this candidate set, we randomly select a smaller subset of hyperparameter tuples and perform pilot training runs for each tuple. (The size of this random subset is reported in each example below.) Pilot training consists of running the training procedures described in Section 3 or Section 4 for a small number of epochs. The tuple of hyperparameter values achieving the best validation loss after pilot training is then selected for full training. Furthermore, we assess robustness by selecting the best $n$ (rather than single best) hyperparameter tuple after pilot training and performing full training with all $n$ of these setups; then we report the best, median, and worst values of the test loss (as in Table 4). We adopt this procedure as it represents a “middle ground” between very refined hyperparameter tuning and completely random selection.
Table 2: Hyperparameter selection, Section 5. Hyperparameter search spaces for PCP-Map and COT-Flow. Here $m$ denotes the size of the context feature or observation ${\bf y}$.
@@
Context Layer width, $u$$\left\{\dfrac{w}{2^{i}}\bigg{|}\dfrac{w}{2^{i}}>m,i=0,1,...\right\}\cup\{m\}$
@@ 6 Numerical Experiments
We test the accuracy, robustness, efficiency, and scalability of our approaches from Sections 3 and 4 using three problem settings that lead to different challenges and benchmark methods for comparison. All numerical experiments are conducted on one Quadro RTX 8000 GPU with 48 GB of RAM. In Section 6.1, we compare our proposed approaches to the Adaptive Transport Maps (ATM) approach developed in [3] on estimating the joint and conditional distributions of six UCI tabular datasets. In Section 6.2, we compare our approaches to a provably convergent approximate Bayesian computation (ABC) approach on accuracy and computational cost using the stochastic Lotka–Volterra model, which yields an intractable likelihood. Using this dataset, we also compare PCP-Map’s and COT-Flow’s sampling efficiency for different settings. In Section 6.3, we demonstrate the scalability of our approaches to higher-dimensional problems by comparing them to the flow-based neural posterior estimation (NPE) approach on an inference problem involving the 1D shallow water equations. To demonstrate the improvements of PCP-Map over the amortized CP-Flow approach in the repository associated with  [25], we compare computational cost in Section 6.4.
@@ 6.1 UCI Tabular Datasets
(30) $J_{{\rm NLL}}[T^{-1}]={\rm\mathbb{E}}_{\pi({\bf x},{\bf y})}\left[{\frac{1}{2}}\left\|T^{-1}({\bf x},{\bf y})\right\|^{2}-\log\det\nabla T^{-1}({\bf x},{\bf y})\right].$
@@
(31) $J_{{\rm NLL},{\bf y}}[\mathsf{h}^{-1}]=\mathbb{E}_{{\bf y}\sim\pi({\bf y})}\left[\left\|\mathsf{h}^{-1}({\bf y})\right\|^{2}-\log\det\nabla_{\bf y}\mathsf{h}^{-1}({\bf y})\right].$
@@
for $k=0,\ldots,K-1$. Here $\sigma^{(s)}$ is the softplus activation function. Since the FICNN map is not part of our contribution, we use the input-augmented ICNN implementation used in [25]. For the second map component $g^{-1}$, we use the gradient of a PICNN as described in Section 3. For COT-Flow, we construct two distinct neural network parameterized potentials $\Phi_{x}$ and $\Phi_{y}$. Here, $\Phi_{y}$ only takes the conditioning variable ${\bf y}$ as inputs and can be constructed exactly like in [41]. $\Phi_{x}$ acts on all features and is the same potential described in Section 4. We learn the inverse conditional generator $g^{-1}_{\theta}$ to perform conditional density estimation. This is equivalent to learning only the weights $\theta$ of the PICNN $\tilde{G}_{\theta}$ for PCP-Map. For COT-Flow, we only construct and learn the potential $\Phi_{x}$.
Table 3: UCI tabular datasets experiment, Section 6.1. Hyperparameter search space.
@@
The hyperparameter search space we use for this experiment is presented in Table 3. We select smaller batch sizes and large learning rates as this leads to fast convergence on these relatively simple problems. For each dataset, we select the ten best hyperparameter combinations based on a pilot run for full training and report the best, median, and worst results. For PCP-Map’s pilot run, we performed 15 epochs for all three conditional datasets, three epochs for the Parkinson’s and the White Wine datasets, and four epochs for Red Wine using 100 randomly sampled combinations. For COT-Flow, we limit the pilot runs to only 50 sampled combinations due to a narrower sample space on model architecture. To assess the robustness of our approaches, we performed five full training runs with random initializations of network weights for each of the ten hyperparameter combinations for each dataset.
In Table 4, we present the best, median, and worst test-time mean NLL as well as the maximum mean discrepancy (MMD) between generated and true samples. For two sample sets $P=\{{\bf p}_{i}\}^{N}_{i=1}$ and $Q=\{{\bf q}_{i}\}^{M}_{i=1}$, the MMD is computed as:
$\text{MMD}(P,Q)=$ $\frac{1}{N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}k({\bf p}_{i},{\bf p}_{j})+\frac{1}{M^{2}}\sum_{i=1}^{M}\sum_{j=1}^{M}k({\bf q}_{i},{\bf q}_{j})$
$-\frac{2}{NM}\sum_{i=1}^{N}\sum_{j=1}^{M}k({\bf p}_{i},{\bf q}_{j}),$
with $k$ chosen to be a squared exponential kernel with bandwidth one, following [23, 48]. Results are reported with empirical standard deviations obtained by repeating training with five random neural network weight initializations (for each fixed set of hyperparameter values). We run our two proposed approaches on six UCI datasets, and compare to the best reported NLL for ATM. The table demonstrates that the best models outperform ATM for all datasets and that the median performance is typically superior. Overall, COT-Flow slightly outperforms PCP-Map in terms of NLL, while the opposite is true in terms of MMD for the best models. The improvements are more pronounced for the conditional density estimation tasks, where even the worst hyperparameters from PCP-Map and COT-Flow improve over ATM by a substantial margin.
Table 4: UCI tabular datasets experiment, Section 6.1. Mean NLL and MMD comparisons on test data, for ATM, PCP-Map, and COT-Flow. For each approach, we report the best, median, and worst results across five training runs and various hyperparameter combinations, with standard deviations shown ($\pm$). Lower values are better, and the best results are highlighted in bold.
@@
Maximum mean discrepancy (MMD) ×$\times 10^{-2}$
PCP-Map (best) 3.3$\pm$0.3 2.8$\pm$0.1 4.3$\pm$0.3 2.7$\pm$1.1 1.4$\pm$0.4 0.7$\pm$0.1
PCP-Map (median) 3.5$\pm$0.4 3.1$\pm$0.3 4.7$\pm$0.3 4.0$\pm$1.1 1.9$\pm$0.5 1.2$\pm$0.5
PCP-Map (worst) 4.0$\pm$0.6 3.3$\pm$0.2 5.5$\pm$0.3 4.8$\pm$1.0 3.8$\pm$1.7 1.7$\pm$0.6
COT-Flow (best) 3.4$\pm$0.3 3.0$\pm$0.1 4.2$\pm$0.1 2.9$\pm$1.4 1.5$\pm$0.3 1.0$\pm$0.3
COT-Flow (median) 4.1$\pm$0.4 3.2$\pm$0.1 4.6$\pm$0.3 4.3$\pm$0.9 3.1$\pm$1.4 1.6$\pm$0.7
COT-Flow (worst) 4.8$\pm$0.2 4.4$\pm$0.1 6.7$\pm$0.3 7.1$\pm$3.0 7.4$\pm$1.9 2.7$\pm$0.3
6.2 Stochastic Lotka–Volterra
We compare our approaches to an ABC approach based on sequential Monte Carlo (SMC) for likelihood-free Bayesian inference using the stochastic Lotka–Volterra (LV) model [63]. The LV model is a stochastic process whose dynamics describe the evolution of the populations $S(t)=(S_{1}(t),S_{2}(t))$ of two interacting species, e.g., predators and prey. These populations start from a fixed initial condition $S(0)=(50,100)$. The parameter ${\bf x}\in\mathbb{R}^{4}$ determines the rate of change of the populations over time, and the observation ${\bf y}\in\mathbb{R}^{9}$ contains summary statistics of the time series generated by the model. This results in an observation vector with nine entries: the mean, the log-variance, the auto-correlation with lags one and two, and the cross-correlation coefficient. The procedure for sampling a trajectory of the species populations is known as Gillespie’s algorithm.
@@ 6.2 Stochastic Lotka–Volterra
Figure 2: Stochastic Lotka–Volterra experiment, Section 6.2. Comparison of estimated posterior distributions and MAP points between the proposed approaches and SMC-ABC (${\bf x}^{*}=(0.01,0.5,1,0.01)^{\top}$). Red dots/bars represent ${\bf x}^{\ast}$; black crosses/bars represent MAP points. Column 1: Proposed approaches (50k training samples). Column 2: Proposed approaches (500k training samples). Column 3: SMC-ABC (around 17.9 million simulations).
@@
We evaluate our approaches for maximum-a-posteriori (MAP) estimation and posterior sampling. We consider a true parameter ${\bf x}^{*}=(0.01,0.5,1,0.01)^{\top}$, which was carefully chosen to give rise to oscillatory behavior in the population time series. Given one observation ${\bf y}^{*}\sim\pi({\bf y}|{\bf x}^{*})$, we first identify the MAP point by maximizing the estimated log-likelihoods provided by our approaches. Then, we generate 2000 samples from the approximate posterior $\pi({\bf x}|{\bf y}^{*})$ using our approaches. Figure 2 presents one and two-dimensional marginal histograms and scatter plots of the MAP point and samples, compared against 2000 samples generated by the SMC-ABC algorithm from [5]. Our approaches yield samples tightly concentrated around the MAP points that are close to the true parameter ${\bf x}^{*}$.
To provide more evidence that our learning approaches indeed solve the amortized problem, Figure 3 shows the MAP points and approximate posterior samples generated from a new random observation ${\bf y}^{*}$ corresponding to the true parameter ${\bf x}^{*}=(0.02,0.02,0.02,0.02)^{\top}$. We observe similar concentrations of the MAP point and posterior samples around the true parameter and similar correlations learned by the generative model and ABC, for example, between the third and other parameters.
@@
Figure 3: Stochastic Lotka–Volterra experiment, Section 6.2. Comparison of estimated posterior distributions and MAP points between the proposed approaches and SMC-ABC (${\bf x}^{*}=(0.02,0.02,0.02,0.02)^{\top}$). Red dots/bars represent ${\bf x}^{\ast}$; black crosses/bars represent MAP points. Column 1: Proposed approaches (50k training samples). Column 2: Proposed approaches (500k training samples). Column 3: SMC-ABC (around 5.8 million simulations).
In the above experiments, we used $n_{t}=32$ to generate posterior samples during testing for the COT-Flow. One can also decrease $n_{t}$ to generate samples faster without sacrificing much accuracy, as shown in Figure 4 and Table 5. To establish such a comparison, we increase the $l$-BFGS tolerance when sampling using PCP-Map, which produces a similar effect on sampling accuracy than decreasing the number of time steps for COT-Flow.
Figure 4: Stochastic Lotka–Volterra experiment, Section 6.2. Relative normed errors between COT-Flow posterior samples generated with smaller number of time-steps $n_{t}$ and samples generated with $n_{t}=32$. Here ${\bf x}^{*}_{1}=(0.01,0.5,1,0.01)^{\top}$ and ${\bf x}^{*}_{2}=(0.02,0.02,0.02,0.02)^{\top}$.
Table 5: Stochastic Lotka–Volterra experiment, Section 6.2. Sampling efficiency comparisons between PCP-Map and COT-Flow in terms of GPU time in seconds (s). We report the mean and standard deviation over five runs, respectively.
@@
The shallow water equations model wave propagation through shallow basins described by the depth profiles parameter, ${\bf x}\in\mathbb{R}^{100}$, which is discretized at 100 equidistant points in space. After solving the equations over a time grid with 100 cells, the resulting wave amplitudes form the 10k-dimensional raw observations. As in [50], we perform a 2D Fourier transform on the raw observations and concatenate the real and imaginary parts since the waves are periodic. We define this simulation-then-Fourier-transform process as our forward model and denote it as $\Psi({\bf x})$. Additive Gaussian noise is introduced into the output of $\Psi$, giving us the observations ${\bf y}=\Psi({\bf x})+0.25{\boldsymbol{\epsilon}}$, where ${\bf y},{\boldsymbol{\epsilon}}\in\mathbb{R}^{200\times 100}$ and ${\boldsymbol{\epsilon}}_{i,j}\sim\mathcal{N}(0,1)$. We aim to use the proposed approaches to learn the posterior $\pi({\bf x}|{\bf y})$.
@@
Table 6: Shallow water experiment, Section 6.3. Hyperparameter search space for the 1D shallow water experiment.
@@
$[\log(\alpha_{1}),\log(\alpha_{2})]$ [$\mathcal{U}(2,5)$, $\mathcal{U}(2,5)$]